# MagicTime: Time-lapse Video Generation Models as Metamorphic Simulators

### v1

@@ Abstract
Recent advances in Text-to-Video generation (T2V) have achieved remarkable success in synthesizing high-quality general videos from textual descriptions. A largely overlooked problem in T2V is that existing models have not adequately encoded physical knowledge of the real world, thus generated videos tend to have limited motion and poor variations. In this paper, we propose MagicTime, a metamorphic time-lapse video generation model, which learns real-world physics knowledge from time-lapse videos and implements metamorphic generation. First, we design a MagicAdapter scheme to decouple spatial and temporal training, encode more physical knowledge from metamorphic videos, and transform pre-trained T2V models to generate metamorphic videos. Second, we introduce a Dynamic Frames Extraction strategy to adapt to metamorphic time-lapse videos, which have a wider variation range and cover dramatic object metamorphic processes, thus embodying more physical knowledge than general videos. Finally, we introduce a Magic Text-Encoder to improve the understanding of metamorphic video prompts. Furthermore, we create a time-lapse video-text dataset called ChronoMagic, specifically curated to unlock the metamorphic video generation ability. Extensive experiments demonstrate the superiority and effectiveness of MagicTime for generating high-quality and dynamic metamorphic videos, suggesting time-lapse video generation is a promising path toward building metamorphic simulators of the physical world.  [^1] Our project aims to help reproduce Sora, supporting Open-Sora-Plan or other DiT-based T2V models.
Figure 1: The illustration of the difference between (a) general videos, and (b) metamorphic videos. Compared to general videos, metamorphic videos contain physical knowledge, long persistence, and strong variation, making them difficult to generate. The general videos are generated by the Animatediff (animatediff, ). Our MagicTime greatly improves capabilities of metamorphic video generation.
@@ 1 Introduction
Recent advances in Text-to-Video (T2V) generation models have been driven by the emergence of diffusion models. These models are typically based on either Transformer architectures (Videopoet, ; Latte, ) or U-Net architectures (VideoComposer, ; Modelscope, ; Align_your_latents, ; animatediff, ). The former approach tokenizes videos and incorporates positional encoding for training, whereas the latter extends the 2D U-Net with a temporal feature extraction module for pseudo-3D training. However, the videos generated by these models primarily exhibit camera movement, lacking a continuous subject metamorphic process and limited incorporation of physical knowledge. We categorize such videos as general videos, illustrated in Fig. 1a. Since the training datasets predominantly consist of general videos, current T2V models struggle to produce videos depicting complex phenomena like seed germination, ice melting, or flower blooming, which are characterized by minimal physical content, brief duration, and limited variation.
Compared to general videos, we have observed a category of videos that typically encompasses the subject’s entire transformation process, thus addressing the inherent limitations of the former. We term this category as metamorphic videos, which encode a more comprehensive representation of world knowledge, as illustrated in Fig. 1b. Such videos have the potential to enhance the capacity of video generation models to accurately depict the real world. Consequently, the generation of metamorphic videos is of paramount importance for a wide range of applications, despite being relatively underexplored in existing research. While the study (freebloom, ) demonstrates the ability to create similar effects by repeatedly inferring and concatenating general videos, it falls short in encoding physical knowledge and lacks generalization through a prompt strategy. In contrast, we aim to develop an end-to-end time-lapse video generation method for adaptively encoding physical knowledge with good generalization.
@@
To tackle the challenge of generating metamorphic videos, we introduce a novel framework, MagicTime, which excels in creating video content that is Magically compressed in Time. Initially, we present the MagicAdapter to encode more physical knowledge in metamorphic videos into the feature extraction, expanding the metamorphic video generation capabilities of pre-trained T2V models without altering their core general video generation abilities. Subsequently, we propose a Dynamic Frames Extraction strategy that enables the model to adapt to the characteristics of time-lapse training videos and prioritize metamorphic video features. Moreover, we introduce a Meta Text-Encoder to refine prompt understanding and distinguish between metamorphic and general prompts. As shown in Fig. 1b, our MagicTime generates text-aligned, coherent, high-quality metamorphic videos across various styles, coordinated with carefully designed techniques.
Table 1: Comparison of the characteristics of our ChronoMagic benchmark with existing time-lapse datasets. We compare time-lapse video characteristics in terms of physics, persistence, and variation.
Dataset Type Strong Physics High Persistence Variation
Sky Time-Lapse ${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[ACMMM'21]}}}$ (SkyTimelapse, ) General
Time-Lapse-D ${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[CVPR'18]}}}$ (TimeLapse-D, ) General
ChronoMagic Metamorphic
Further, we meticulously collect a dataset, ChronoMagic, sourced from Internet time-lapse videos for training our method. As shown in Table 1, our ChronoMagic dataset is distinct from previous time-lapse datasets (SkyTimelapse, ; TimeLapse-D, ), as it consists of metamorphic time-lapse videos (e.g. ice melting, and flower blooming) that show strong physics, high persistence, and variation. Given the domain discrepancy between metamorphic and general videos, we propose the integration of Cascade Preprocessing and Multi-View Text Fusion to capture dynamics into detailed video captions. Contributions are summarized as follows:
  • We introduce MagicTime, a metamorphic time-lapse video generation diffusion model by adopting a standard T2V model with a MagicAdapter scheme. This addresses the limitations of existing methods that are unable to generate metamorphic videos.
@@
  • We propose an automatic metamorphic video captioning annotation framework. Utilizing this framework, we have curated a high-quality dataset named ChronoMigic, consisting of 2,265 time-lapse videos, each accompanied by a detailed caption.
@@ 2 Related Work
Text-to-Image Generation    Text-to-image (T2I) synthesis has made significant strides thanks to advancements in deep learning and the availability of large image-text datasets such as CC3M (CC3M, ) and LAION (LAION, ). Compared to variational autoencoders (VAEs) (VAE, ; vqVAE, ), flow-based models (Rezende_Mohamed_2015, ; Dinh_Sohl-Dickstein_Bengio_2016, ), and generative adversarial networks (GANs) (tedigan, ; StyleGAN-XL, ; StyleGAN-T, ), diffusion models (Imagen, ; DALLE, ; DALLE2, ; sdxl, ; lian2023llm, ; Multidiffusion, ) have demonstrated exceptional efficacy in image modeling. For instance, contemporary T2I models often employ Latent Diffusion Models (LDM) (LDM, ), which utilize sampling in latent spaces from pre-trained autoencoders to minimize computational overhead. Recapti (Recapti, ) harnessed the power of Large Language Models (LLM) to improve the text comprehension capabilities of LDM, enabling training-free T2I generation. Furthermore, the integration of additional control signals (ucontrolnet, ; chen2024training, ; t2iadapter, ; dragdiffusion, ; Gligen, ) has further expanded the applicability of Diffusion Models in T2I synthesis. Unlike image generation, the task of video generation required by MagicTime is more complex, as it necessitates both diversity in content across space and consistency in content across time.
@@
Figure 2: The pipeline for high-quality time-lapse video data annotation. (a) Cascade Preprocessing. We successfully detect many useless cuts. The results are obtained by OpenCV. (b) Multi-view Text Fusion. Compared with only inputting keyframes, our pipeline can obtain vivid, detailed and accurate descriptions.
@@
Text-to-Video Generation    The advent of large-scale video-text datasets (webvid, ; Internvid, ; HD-VG-130M, ) has facilitated significant advancements in pre-trained video diffusion models (Make-a-video, ; SVD, ; Align_your_latents, ; ge2023preserve, ; Modelscope, ; Show-1, ). MagicVideo2 (MagicVideo-V2, ) utilized cascaded diffusion to separate image and video generation, starting with text to initially generate images and subsequently produce high-quality videos. Videopoet (Videopoet, ) employed a Transformer-based diffusion approach, combined with multimodal inputs, for zero-shot video generation. Animatediff (animatediff, ) and Lumiere (lumiere, ) extended the 2D-UNet to a 3D-UNet by adding temporal layers and implementing a latent denoising process to conserve computational resources. While these models have achieved basic video generation, the resulting videos typically feature a single scene and a single action, and lack encoding of physical world knowledge. In contrast, our MagicTime incorporates more physics knowledge and offers the capability to generate metamorphic videos.
@@
Time-lapse Video Generation    Previous studies (nam2019end, ; FGLA, ; DtvNet, ; AL, ) on time-lapse video generation predominantly concentrated on general videos, failing to produce the metamorphic videos that we address. For example, (FGLA, ; DtvNet, ) treated time-lapse videos merely as a distinct domain, overlooking the physical knowledge embedded within them. Consequently, the generated videos were limited to scenes with relatively minor movements, such as drifting clouds or moving sunsets. In contrast, our work primarily focuses on generating metamorphic time-lapse videos, which exhibit larger variations and depict the entire process of object metamorphosis. In this work, we aim to incorporate more physical knowledge into MagicTime for metamorphic time-lapse video generation.
@@ 3 Methodology
In this section, we commence with a brief overview of the latent diffusion model. Subsequently, we describe the construction of the ChronoMagic dataset. Lastly, we present an outline of MagicTime (refer to Fig. 3) and elaborate on its implementation details.
@@ 3.1 Preliminaries
Diffusion models (DDPM, ; DDIM, ) typically consist of two processes: forward diffusion and reverse denoising. Given an input signal $x_{0}$, the process of forward diffusion is defined as follows:
@@
where $t=1,...,T$ are time steps, and $\beta_{t}\in(0,1)$ is the noise schedule. When the total time steps $T$ are sufficiently large, the resulting $x_{t}$ eventually approximates Gaussian noise $\in N(0,I)$. The purpose of the denoising process is to learn how to reverse the forward diffusion and gradually remove the noise as follows:
@@
$\mathcal{L}=\mathbb{E}_{,y,\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,\tau_{\theta}(y)\right)\right\|_{2}^{2}\right],$ (3)
where $y$ is text condition, and $\tau_{\theta}(\cdot)$ is the text encoder. By replacing $x_{0}$ with $\mathcal{E}\left(x_{0}\right)$, the latent diffusion model is derived, which also serves as the loss function for MagicTime. In video diffusion models, the parameter $\theta$ typically employs either a Transformer (Videopoet, ; Latte, ) or a U-Net architecture (VideoComposer, ; Modelscope, ; Align_your_latents, ; animatediff, ; zhang2024continuous, ). Considering that the most effective open-source T2V models currently utilize U-Net architecture, we have embraced this architecture in this work.
Figure 3: Overview of the proposed MagicTime approach. We first train MagicAdapter-S to remove the influence of watermarks. Next, MagicAdapter-T is trained to generate metamorphic videos with the help of Dynamic Frames Extraction. Finally, we train a Magic Text-Encoder to enhance text comprehension ability. During the inference stage, all components need to be loaded simultaneously. Slash padding indicates the module is not used.
@@ 3.2 ChronoMagic Dataset
Data Curation and Filter    We found that videos in large-scale text-video datasets like WebVid (webvid, ) typically consist of coherent segments depicting everyday life (general videos), devoid of significant changes or object metamorphosis. Consequently, open-source models (SVD, ; Make-a-video, ; animatediff, ) are limited to generating general videos and are unable to produce metamorphic videos. To overcome this limitation, we introduce the ChronoMagic dataset, which includes 2,265 high-quality time-lapse videos with accompanying text descriptions from open-domain content. Time-lapse videos capture complete object metamorphosis and convey more physical knowledge than general videos, thereby enhancing generative models’ understanding of the world.
Given that video platforms host a multitude of low-quality videos, indiscriminate collection would inevitably compromise the dataset’s quality. To mitigate this, we leverage the metadata from video websites for collection and filtering. Our data collection process begins with retrieving original videos from YouTube using "time-lapse" as a search term. Subsequently, we implement a filtering mechanism that excludes videos with short titles, low view counts, or absent hashtags. Furthermore, we discard videos with irrelevant hashtags such as "YouTube," "video," and "shorts." Building on the annotations generated in Section 3.2, we further refine our selection by eliminating irrelevant videos, thus establishing a closed-loop data collection and iteration process. Ultimately, we curate a collection of 2,265 time-lapse videos that meet our specified criteria.
Cascade Preprocessing    Due to complex scene transitions in original videos (see Fig. 2a) affecting the model’s learning of temporal relevance, we need to find a way to solve this problem. The most direct method is to use existing tools for detection, so we choose OpenCV (opencv, ) as a tool. In the initial stage, we define $F$ as the set of all video frames. Each frame $f_{i}$ in $F$ is converted to grayscale $g_{i}\in(1,H,W)$ to mitigate the influence of color on detection. Subsequently, OpenCV is used to calculate the pixel intensity difference $\mathcal{D}_{i,i+1}$ between consecutive frames, yielding the average pixel intensity $\mathcal{S}_{i,i+1}^{opencv}$:
@@
$\mathcal{S}_{i,i+1}^{opencv}=\sum_{h=1}^{H}\sum_{w=1}^{W}\mathcal{D}_{i,i+1}(h,w),\quad\mathrm{for~{}}i=1\mathrm{~{}to~{}}F-1$ (5)
if $\mathcal{S}_{i,i+1}^{opencv}>\theta$, then this position is defined as the transition point. However, we noted that OpenCV is prone to occasional errors. Consequently, we further employ CLIP (CLIP, ) to detect differences between frames:
@@
$\mathcal{T}_{i}=\begin{cases}\text{True,}&\text{if }\mathcal{S}_{i,i+1}^{opencv}>\theta\text{ and }\mathcal{S}_{i,i+1}^{clip}<\vartheta\\ \text{False,}&\text{otherwise}\end{cases}$ (7)

### v2

@@ Abstract
Recent advances in text-to-video generation (T2V) have achieved remarkable success in synthesizing high-quality general videos from textual descriptions. A largely overlooked problem in T2V is that existing models have not adequately encoded physical knowledge of the real world, thus generated videos tend to have limited motion and poor variations. In this paper, we propose MagicTime, a metamorphic time-lapse video generation model, which learns real-world physics knowledge from time-lapse videos and implements metamorphic generation. First, we design a simple yet effective two-stage Magic Adaptive Strategy, encode more physical knowledge from metamorphic videos, and transform pre-trained T2V models to generate metamorphic videos. Second, we introduce a Dynamic Frames Extraction strategy to adapt to metamorphic time-lapse videos, which have a wider variation range and cover dramatic object metamorphic processes, thus embodying more physical knowledge than general videos. Finally, we introduce a Magic Text-Encoder to improve the understanding of metamorphic video prompts. Furthermore, we create a time-lapse video-text dataset called ChronoMagic, specifically curated to unlock the metamorphic video generation ability. Extensive experiments demonstrate the superiority and effectiveness of MagicTime for generating high-quality and dynamic metamorphic videos, suggesting time-lapse video generation is a promising path toward building metamorphic simulators of the physical world.
Index Terms:
Text-to-Video Generation, Diffusion Model, Time-lapse Video, Metamorphic Video
Figure 1: The illustration of the difference between (a) general videos, and (b) metamorphic videos. Compared to general videos, metamorphic videos contain physical knowledge, long persistence, and strong variation, making them difficult to generate. The general videos are generated by the Animatediff [1]. The metamorphic videos generated by our MagicTime demonstrate that it can greatly improve the capabilities of metamorphic video generation.
@@ 1 Introduction
Recent advances in text-to-video (T2V) generation models have been driven by the emergence of diffusion models. These models are typically based on either Transformer architectures [2, 3] or U-Net architectures [4, 5, 6, 1]. The former approach tokenizes videos and incorporates positional encoding for training, whereas the latter extends the 2D U-Net with a temporal feature extraction module for pseudo-3D training. However, the videos generated by these models primarily exhibit camera movement, lacking a continuous subject metamorphic process and limited incorporation of physical knowledge. We categorize such videos as general videos, illustrated in Fig. 1a. Since the training datasets predominantly consist of general videos, current T2V models struggle to produce videos depicting complex phenomena like seed germination, ice melting, or flower blooming, which are characterized by minimal physical content, brief duration, and limited variation.
TABLE I: Comparison of the characteristics of our ChronoMagic benchmark with existing time-lapse datasets. We compare time-lapse video characteristics in terms of physics, persistence, and variation.
Dataset Strong Physics High Persistence Variation
Sky Time-Lapse [7]
Time-Lapse-D [8]
ChronoMagic
Compared to general videos, we have observed a category of videos that typically encompasses the subject’s entire transformation process, thus addressing the inherent limitations of the former. We term this category as metamorphic videos, which encode a more comprehensive representation of world knowledge, as illustrated in Fig. 1b. Such videos have the potential to enhance the capacity of video generation models to accurately depict the real world. Consequently, the generation of metamorphic videos is of paramount importance for a wide range of applications, despite being relatively underexplored in existing research. While the study [9] demonstrates the ability to create similar effects by repeatedly inferring and concatenating general videos, it falls short in encoding physical knowledge and lacks generalization through a prompt strategy. In contrast, we aim to develop an end-to-end time-lapse video generation method for adaptively encoding physical knowledge with strong generalization capabilities.
@@
To tackle the challenge of generating metamorphic videos, we introduce a novel framework, MagicTime, which excels in creating video content that is Magically compressed in Time. Initially, we present the Magic Adaptive Strategy to encode more physical knowledge in metamorphic videos into the feature extraction, expanding the metamorphic video generation capabilities of the pre-trained T2V model. Subsequently, we propose a Dynamic Frames Extraction strategy that enables the model to adapt to the characteristics of time-lapse training videos and prioritize metamorphic video features without altering its core general video generation abilities. Moreover, we introduce a Meta Text-Encoder to refine prompt understanding and distinguish between metamorphic and general prompts. As shown in Fig. 1b, our MagicTime generates text-aligned, coherent, high-quality metamorphic videos across various styles, coordinated with carefully designed techniques.
Further, we meticulously collect a dataset, ChronoMagic, sourced from Internet time-lapse videos for training our method. As shown in Table I, our ChronoMagic dataset is distinct from previous time-lapse datasets [7, 8], as it consists of metamorphic time-lapse videos (e.g. ice melting, and flower blooming) that show strong physics, high persistence, and variation. Given the domain discrepancy between metamorphic and general videos, we propose the integration of Cascade Preprocessing and Multi-View Text Fusion to capture dynamics into detailed video captions. Contributions are summarized as follows:
  • We introduce MagicTime, a metamorphic time-lapse video generation diffusion model, by adopting a standard T2V model with a Magic Adaptive Strategy. This addresses the limitations of existing methods that are unable to generate metamorphic videos.
@@
  • We propose an automatic metamorphic video captioning annotation framework. Utilizing this framework, we have curated a high-quality dataset named ChronoMagic, consisting of 2,265 time-lapse videos, each accompanied by a detailed caption.
@@ 2 Related Work
Text-to-Image Generation    Text-to-image (T2I) synthesis has made significant strides thanks to advancements in deep learning and the availability of large image-text datasets such as CC3M [10] and LAION [11]. Compared to variational autoencoders (VAEs) [12, 13], flow-based models [14, 15], and generative adversarial networks (GANs) [16, 17, 18], diffusion models [19, 20, 21, 22, 23, 24] have demonstrated exceptional efficacy in image modeling. For instance, contemporary T2I models often employ Latent Diffusion Models (LDM) [25], which utilize sampling in latent spaces from pre-trained autoencoders to minimize computational overhead. Recapti [26] harnessed the power of Large Language Models (LLM) to improve the text comprehension capabilities of LDM, enabling training-free T2I generation. Furthermore, the integration of additional control signals [27, 28, 29, 30, 31] has further expanded the applicability of Diffusion Models in T2I generation. Unlike image generation, the task of video generation required by MagicTime is more complex, as it necessitates both diversity in content across space and consistency in content across time.
@@
Figure 2: The pipeline for high-quality time-lapse video data annotation. (a) Cascade Preprocessing. It automatically divides a video into clips at corresponding transitions in each stage, thus dividing a complex raw video into clips. N-clips/video refers that a video needs to be divided into N clips, and the smaller the value, the more coherent the video. (b) Multi-view Text Fusion. Compared with only inputting keyframes, our pipeline can obtain vivid, detailed, and accurate descriptions.
@@
Text-to-Video Generation    The advent of large-scale video-text datasets [32, 33, 34] has facilitated significant advancements in pre-trained video diffusion models [35, 36, 37, 6, 38, 5, 39, 40, 41, 42]. MagicVideo2 [43] utilized cascaded diffusion to separate image and video generation, starting with text to initially generate images and subsequently produce high-quality videos. Videopoet [2] employed a Transformer-based diffusion approach, combined with multimodal inputs, for zero-shot video generation. Animatediff [1], and Lumiere [44] extended the 2D-UNet to a 3D-UNet by adding temporal layers and implementing a latent denoising process to conserve computational resources. While these models have achieved basic video generation, the resulting videos typically feature a single scene and a single action, and lack encoding of physical world knowledge. In contrast, our MagicTime incorporates more physics knowledge and offers the capability to generate metamorphic videos.
@@
Time-lapse Video Generation    Previous studies [45, 46, 47, 48] on time-lapse video generation predominantly concentrated on general videos, failing to produce the metamorphic videos that we address. For example, [46, 47] treated time-lapse videos merely as a distinct domain, overlooking the physical knowledge embedded within them. Consequently, the generated videos were limited to scenes with relatively minor movements, such as drifting clouds or moving sunsets. In contrast, we primarily focus on generating metamorphic time-lapse videos, which exhibit significant variations and depict the complete process of object metamorphosis. The goal of our work is to integrate more physical knowledge into MagicTime for the generation of high-quality metamorphic time-lapse videos.
@@ 3 Methodology
This section begins with a brief overview of the diffusion model, followed by a description of the ChronoMagic dataset construction. Finally, we provide an outline of MagicTime (see Fig. 3) and discuss its implementation details.
@@ 3.1 Preliminaries
Diffusion models [49, 50, 51] typically consist of two processes: forward diffusion and reverse denoising. Given an input signal $x_{t}$, the process of forward diffusion is defined:
@@
where $t=1,...,T$ are time steps, and $\beta_{t}\in(0,1)$ is the noise schedule. When the total time steps $T$ are sufficiently large, the resulting $x_{t}$ eventually approximates Gaussian noise $\in N(0,I)$. The purpose of the denoising process is to learn how to reverse the forward diffusion:
@@
$\mathcal{L}=\mathbb{E}_{x_{t},y,\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},t,\tau_{\theta}(y)\right)\right\|_{2}^{2}\right],$ (3)
where $y$ is text condition, and $\tau_{\theta}(\cdot)$ is the text encoder. By replacing $x_{t}$ with $\mathcal{E}\left(x_{t}\right)$, the latent diffusion model is derived [25, 52], which also serves as the loss function for MagicTime. In video diffusion models, the parameter $\theta$ typically employs either a Transformer [2, 3] or a U-Net architecture [53, 54, 55, 56, 4, 5, 6, 1]. Considering that the most effective open-source T2V models currently utilize U-Net, we have embraced this architecture in this work.
Figure 3: Overview of the proposed MagicTime approach. We first train MagicAdapter-S to focus on static state information. Next, MagicAdapter-T is trained to encode dynamic stage change with the help of Dynamic Frames Extraction. Finally, we train a Magic Text-Encoder to enhance text comprehension ability. During the inference stage, all components need to be loaded simultaneously. Slash padding indicates the module is not used.
@@ 3.2 ChronoMagic Dataset
Data Curation and Filter    We found that videos in existing large-scale text-video datasets like WebVid [32] typically consist of coherent segments depicting everyday life (general videos), devoid of significant changes or object metamorphosis. Consequently, open-source models [37, 36, 1] are limited to generating general videos and are unable to produce metamorphic videos. To overcome this limitation, we introduce the ChronoMagic dataset, which includes 2,265 high-quality time-lapse videos with accompanying text descriptions from open-domain content. Time-lapse videos capture complete object metamorphosis and convey more physical knowledge than general videos, thereby enhancing generative models’ understanding of the world.
Given that video platforms host a multitude of low-quality videos, indiscriminate collection would inevitably compromise the dataset’s quality. To mitigate this, we leverage the metadata from video websites for collection and filtering. Our data collection process begins with retrieving original videos from YouTube using ”time-lapse” as a search term. Subsequently, we implement a filtering mechanism that excludes videos with short titles, low view counts, or absent hashtags. Furthermore, we discard videos with irrelevant hashtags such as ”YouTube,” ”video,” and ”shorts.” Building on the annotations generated in Section 3.2, we further refine our selection by eliminating irrelevant videos, thus establishing a closed-loop data collection and iteration process. Ultimately, we curate a collection of 2,265 time-lapse videos that meet our specified criteria.
Cascade Preprocessing    Due to complex scene transitions in Internet videos (see Fig. 2a) affecting the model’s learning of temporal relevance, we need to find a way to solve this problem. We first define $F$ as the set of all video frames. In the initial stage, each frame $f_{i}$ in $F$ is converted to grayscale $g_{i}\in(1,H,W)$ to mitigate the influence of color on detection. Subsequently, we calculate the pixel intensity difference $\mathcal{D}_{i,i+1}$ between consecutive frames, yielding the average pixel intensity $\mathcal{S}_{i,i+1}$:
@@
$\mathcal{S}_{i,i+1}=\sum_{h=1}^{H}\sum_{w=1}^{W}\mathcal{D}_{i,i+1}(h,w),\quad\mathrm{for~{}}i=1\mathrm{~{}to~{}}F-1$ (5)
if $\mathcal{S}_{i,i+1}>\theta$, then this position is defined as the transition point. However, we noted that this simple method is prone to occasional errors. Consequently, in stage two, we further employ CLIP [57] to detect transitions:
@@
$\mathcal{T}_{i}=\begin{cases}\text{True,}&\text{if }\mathcal{S}_{i,i+1}>\theta\text{ and }\mathcal{S}_{i,i+1}^{clip}<\vartheta\\ \text{False,}&\text{otherwise}\end{cases}$ (7)